# Simple fits for the neutrino luminosities from protoneutron star cooling

### v2

@@ IV.1 Class A EoS
  • The normalization parameter $C$ (top panels) adopts values of $5\,\operatorname{B}/\operatorname{s}\lesssim C\lesssim 15\,\operatorname{B}/\operatorname{s}$ for all neutrino species and it increases with the PNS mass. In particular, DD2 leads to the lowest values and SFHo and SFHx feature a similar dependence, with SFHx providing the largest values of $C$.
@@
  • The power-law index $\alpha$ (second panels from top) shows a mild dependence on the PNS mass, featuring values of $0.50\lesssim\alpha\lesssim 0.70$ for all neutrino species, with an increase in the case of DD2 and a slight decrease for SFHo and SFHx. Similar to the normalization parameter, for each kind of neutrino and PNS mass, the largest value of $\alpha$ is obtained with SFHx and the smallest one with DD2.
@@
  • The late-time suppression parameters $\tau$ and $n$ (third and forth panels from top) tend to increase with the PNS mass for all neutrino species, with values of $5\,\operatorname{s}\lesssim\tau\lesssim 10\,\operatorname{s}$ and $3\lesssim n\lesssim 5$. In particular, for all neutrino kinds and PNS masses, DD2 leads to the lowest values of $\tau$ and SFHx to the largest ones. On the other hand, DD2 yields the largest $n$ values for $\nu$ and $\bar{\nu}$ of all flavors and all PNS masses, whereas the $1\,\sigma$ confidence bands of $n$ for SFHo and SFHx models overlap in the mass range of $1.4\,M_{\odot}\lesssim M_{\rm NS}\lesssim 1.7\,M_{\odot}$, with SFHx showing larger values of $n$ than SFHo at higher PNS masses.
@@
A given value of $C$, $\tau$, and $n$ can be obtained with different combinations of EoS and NS mass. For instance, for $\bar{\nu}_{e}$, $\tau=7\,\operatorname{s}$ corresponds to $M_{\rm NS}\approx 1.38\,M_{\odot}$ and SFHx, $M_{\rm NS}\approx 1.45\,M_{\odot}$ and SFHo, as well as $M_{\rm NS}\approx 1.77\,M_{\odot}$ and DD2. However, the best-fit values of $\alpha$ lie in different ranges for different EoSs. For instance, $0.6\lesssim\alpha_{\bar{\nu}_{e}}\lesssim 0.65$ would point to the SFHo EoS, irrespective of the PNS mass. Therefore, in the case of a future observation of a SN neutrino signal, the measurement of $\alpha$ would provide information on the EoS and the PNS mass and, combined with the measurements of the other fit parameters, would allow us to characterize the PNS mass and the EoS. More explicitly, neglecting for simplicity flavor mixing, a combined measurement of $\tau=7\,\operatorname{s}$, $\alpha\approx 0.63$, and $n\approx 3.9$ for the $\bar{\nu}_{e}$ luminosity would suggest a SN explosion leading to a PNS of mass $1.44~{}M_{\odot}$, whose interior properties are described by the SFHo EoS. As further discussed in Sec. VII, due to the similarities of the spectra and luminosities of electron antineutrinos and heavy-lepton antineutrinos in our models at times $t\gtrsim 1$ s, flavor conversions are not a major effect, although they would make the reconstruction of the fit parameters less straightforward, but without spoiling our result. The accurate reconstruction of the parameters is left for future work.
@@ IV.2 LS220 EoS
  • The normalization parameter $C$ (top panels) is $C\sim O(10-1000)\,{\rm B}/\operatorname{s}$ for all of the neutrino species and it increases linearly with the PNS mass.
  • The power-law index $\alpha$ (second panels from top) tends to decrease as the PNS mass increases. It is negative for all models and neutrino species except for $\nu_{e}$ and $\bar{\nu}_{e}$ in the simulation with $M_{\rm NS}=1.36\,M_{\odot}$, where $\alpha>0$.
  • The late-time suppression parameters $\tau$ and $n$ (third and forth panels from top) tend to decrease with higher PNS mass. For the 1.36-LS220 model we get $\tau\sim 4$ s and $n>1$ for $\nu_{e}$ and $\bar{\nu}_{e}$ (see Tab. 5 in Appendix B), whereas $\tau\lesssim 0.25$ s and $n\lesssim 0.5$ in all the other cases (Tables 6 and 7).
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 11: Correlations of the fitting parameters $\tau$ (in units of seconds) and $\alpha$ for $\nu_{e}$ (top left panel), $\bar{\nu}_{e}$ (top right panel), $\nu_{\mu}$ (central left panel), $\bar{\nu}_{\mu}$ (central right panel), $\nu_{\tau}$ (bottom left panel), $\bar{\nu}_{\tau}$ (bottom right panel), for all Case A EoSs and all considered PNS masses, obtained from simulation output data between 1 s and $t_{\nu_{i},{\rm c}}$ for each neutrino species $\nu_{i}$.
@@ V Correlation between $\tau$ and $\alpha$
for all neutrino species except $\bar{\nu}_{e}$. This is shown in Fig. 11 for neutrinos (left) and antineutrinos (right) of all flavors and all Class A EoSs. In Appendix D, in Table 9 we report the best-fit values with errors obtained from linear regression for $A$ and $B$ for all neutrino species. The pairs of values $(\tau,\alpha)$ are represented by filled circles for the DD2 EoS, filled squares for SFHo, and filled diamonds for SFHx in Fig. 11. Different colours correspond to the different PNS masses: black for $M_{\rm NS}=1.36~{}M_{\odot}$, red for $M_{\rm NS}=1.44~{}M_{\odot}$, blue for $M_{\rm NS}=1.62~{}M_{\odot}$, green for $M_{\rm NS}=1.77~{}M_{\odot}$, and orange for $M_{\rm NS}=1.93~{}M_{\odot}$.
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ VI Impact of muons and convection
In order to assess the impact of some of our physics inputs of the simulations, we investigate additional PNS cooling calculations for a PNS mass of 1.62 $M_{\odot}$ now, where we either omitted convection or muons. Non-convective models were computed for all considered EoSs and are denoted by a suffix “-c” appended to their names (e.g., 1.62-DD2-c), whereas only two models are considered without muons (suffix “-m”), namely 1.62-DD2-m and 1.62-SFHo-m [22]. The bottom two data blocks of Tab. 1 provide the final simulation times and the corresponding reduction factors $X_{\nu_{i}}^{\rm fin}$ for all neutrino species in these additional models. Bold numbers in Tab. 1 for models 1.62-SFHx-c and 1.61-LS220-c signal that the simulations were stopped when $X_{\nu_{i}}^{\rm fin}>0.15$, implying that for these models our standard cut-off time $t_{\nu_{i},{\rm c}}$ is larger than $t_{\rm fin}$. In both of the simulations, the weakest suppression is obtained for $\nu_{\mu}$, with $X_{\nu_{\mu}}^{\rm fin}=0.219$ for 1.62-SFHx-c and $X_{\nu_{\mu}}^{\rm fin}=0.185$ for 1.61-LS220-c (see the values marked by a star in Tab. 1). Therefore, to test the impact of convection for the SFHx and LS220 EoSs we cut our luminosity data at $t_{\nu_{i},{\rm Co}}$, when $X_{\nu_{i}}^{{\rm Co}}=t_{\nu_{i},{\rm Co}}\,L_{\nu_{i},{\rm Co}}/t_{\nu_{i},{\rm max}}\,L_{\nu_{i},{\rm max}}=0.22$ for simulations with SFHx and $X_{\nu_{i}}^{{\rm Co}}=0.19$ for simulations with LS220. The values of $t_{\nu_{i},{\rm Co}}$ for all neutrino species of these simulations are given in Tab. 4 of Appendix A.
@@
In the absence of convection the quantity $t\,L_{\nu}$ declines steeply only at $t\gtrsim 10\,\operatorname{s}$ for all neutrino species, with different characteristic features depending on the EoS (see the second data block from the bottom of Tables 2 and 3 in Appendix A). In particular, we witness that
@@
  • LS220 (red lines) leads to a peak in $t\,L_{\nu}$ at $t\approx 2$ s and shows a more shallow decline afterwards, forming a plateau-like shape in the time interval $2~{}\operatorname{s}\lesssim t\lesssim 10$ s (i.e., $L_{\nu}$ follows approximately $L_{\nu}\propto t^{-1}$) before a steeper decline sets at $t\gtrsim 15$ s.
@@
  • the absence of convection leads to a considerable stretching of the PNS Kelvin-Helmholtz neutrino cooling time, with the most moderate change for DD2,
  • the omission of muons has a relatively mild effect on the evolution of the neutrino signals for the displayed 1.62 $M_{\odot}$ models, making the suppression in the luminosity only slightly faster (because the NS becomes less compact with a lower binding energy), as visible by the green lines in the left panels.
To quantitatively assess the impact of these variations of the input physics of our models, we also fit the neutrino and antineutrino signals of the additional simulations with the expression of Eq. (2) and compare the best-fit parameters with those obtained in our benchmark simulations. As an example, Fig. 14 presents the simulation data (blue) and their best fits (orange) in the time interval of interest for $\bar{\nu}_{e}$ results from the simulations without convection (DD2 and SFHo in the top panels and SFHx and LS220 in the middle panels) and from our two simulations without muons (bottom panels). For the simulations without muons, where the neutrino signal does not experience major changes, the agreement between data and fits is similarly good as for our benchmark simulations. For the simulations without convection the fit is still of excellent quality for DD2, although the omission of convection has altered the shape of the curve of $tL_{\bar{\nu}_{e}}$ (see Fig. 13). In contrast, we obtain visibly larger discrepancies between fits and data for the non-convective simulations with SFHo and SFHx, for which in particular the shape of $tL_{\bar{\nu}_{e}}$ in the power-law dominated early phase cannot be reproduced as well as for models that include convection. Notably, the fits for $\bar{\nu}_{e}$ (and similarly for all other neutrino species) slightly overestimate $tL_{\nu}$ at $t\approx 1\,\operatorname{s}$ and tend to peak only at somewhat later times. Finally, the plateau-like region of $tL_{\nu}$ in the simulation results for LS220 implies a best-fit value of $\alpha\approx 1$, with the luminosity fit overestimating the data at $t\approx 1\,\operatorname{s}$ and following well a $t^{-1}$ power law before being exponentially suppressed at $t\gtrsim 10\,\operatorname{s}$. Our findings for all other neutrino species are analogous. In Tables 10 and 11 of Appendix E we provide the best-fit parameter values and their $1\sigma$ errors for electron and muon neutrinos and antineutrinos, for our $1.62\,M_{\odot}$ models with different EoSs and varied input physics.
@@
which implies that the exponential luminosity decline starts at later times and also proceeds more slowly. For $\bar{\nu}_{e}$ in model 1.62-DD2-c we notice an exception from the described general trends with respect to $\tau_{\bar{\nu}_{e}}$, which is slightly smaller than the value of the corresponding model with convection (see Tab. 10 and Fig. 16 in Appendix E). In this case the mild decrease of $\tau_{\bar{\nu}_{e}}$ seems to be compensated by a reduction of $n_{\bar{\nu}_{e}}$ by a factor $\sim$3 compared to the non-convective model, which is by far the largest relative change for any neutrino species in all models with vs. without convection. Interestingly, the change in $\alpha$ depends on the EoS and neutrino species, showing for instance a decrease in 1.62-DD2-c compared to 1.62-DD2, and an increase in 1.62-SFHo-c compared to 1.62-SFHo, and a decrease or slight increase in 1.62-SFHx-c compared to 1.62-SFHx depending on the type of neutrino. This nonuniform behavior points to differences in the influence of the EoS on PNS convection and the associated effects on the emission of different kinds of neutrinos during the early PNS cooling phase.
For simulations with the LS220 EoS, partly because of the poorer quality of the fits for the benchmark models, the omission of convection leads to radical changes in the values of the best-fit parameters. Indeed, LS220 simulations without convection show positive values of $\alpha$ (around unity), much larger values of $\tau$ and $n$ compared to the full-physics cases, and values of $C$ that are well compatible with those of simulations with the other EoSs including and excluding convection, i.e., the $C$-values are close to the luminosity values at 1 s instead of being several 100 B/s for our benchmark models. Finally, the weaker impact of muons on the neutrino signal is highlighted by the small changes in the best-fit parameters obtained for simulations without muons, as further detailed in Appendix E.
VII Counting rate in future detectors
In order to exemplify a possible application of our luminosity fits, we discuss in this section the time evolution of the counting rate $tR_{\nu}(t)$ in a neutrino detector that will monitor the $tL_{\nu}(t)$ evolution in the case of a future Galactic SN explosion. We will demonstrate that our fitting recipe is also useful for fitting the observed neutrino signal. For this purpose, we consider as a reference case a SN at a distance of $D=10$ kpc and evaluate the predicted signal in the future water Cherenkov detector of HyperKamiokande (HK), inspired by the analysis in Ref. [29].

### v3

@@ IV.1 Class A EoS
    The “normalization parameter $C$” (top panels) adopts values of $5\,\operatorname{B}/\operatorname{s}\lesssim C\lesssim 15\,\operatorname{B}/\operatorname{s}$ for all neutrino species and it increases with the PNS mass. In particular, DD2 leads to the lowest values and SFHo and SFHx feature a similar dependence, with SFHx providing the largest values of $C$.
  2. (ii)
@@
    The “power-law index” $\alpha$ (second panels from top) shows a mild dependence on the PNS mass, featuring values of $0.50\lesssim\alpha\lesssim 0.70$ for all neutrino species, with an increase in the case of DD2 and a slight decrease for SFHo and SFHx. Similar to the normalization parameter, for each kind of neutrino and PNS mass, the largest value of $\alpha$ is obtained with SFHx and the smallest one with DD2.
  3. (iii)
@@
    The “late-time suppression parameters” $\tau$ and $n$ (third and fourth panels from top) tend to increase with the PNS mass for all neutrino species, with values of $5\,\operatorname{s}\lesssim\tau\lesssim 10\,\operatorname{s}$ and $3\lesssim n\lesssim 5$. In particular, for all neutrino kinds and PNS masses, DD2 leads to the lowest values of $\tau$ and SFHx to the largest ones. On the other hand, DD2 yields the largest $n$ values for $\nu$ and $\bar{\nu}$ of all flavors and all PNS masses, whereas the $1\,\sigma$ confidence bands of $n$ for SFHo and SFHx models overlap in the mass range of $1.4\,M_{\odot}\lesssim M_{\rm NS}\lesssim 1.7\,M_{\odot}$, with SFHx showing larger values of $n$ than SFHo at higher PNS masses.
@@
A given value of $C$, $\tau$, and $n$ can be obtained with different combinations of EoS and neutron star (NS) mass. For instance, for $\bar{\nu}_{e}$, $\tau=7\,\operatorname{s}$ corresponds to $M_{\rm NS}\approx 1.38\,M_{\odot}$ and SFHx, $M_{\rm NS}\approx 1.45\,M_{\odot}$ and SFHo, as well as $M_{\rm NS}\approx 1.77\,M_{\odot}$ and DD2. However, the best-fit values of $\alpha$ lie in different ranges for different EoSs. For instance, $0.6\lesssim\alpha_{\bar{\nu}_{e}}\lesssim 0.65$ would point to the SFHo EoS, irrespective of the PNS mass. Therefore, in the case of a future observation of a SN neutrino signal, the measurement of $\alpha$ would provide information on the EoS and the PNS mass and, combined with the measurements of the other fit parameters, would allow us to characterize the PNS mass and the EoS. More explicitly, neglecting for simplicity flavor mixing, a combined measurement of $\tau=7\,\operatorname{s}$, $\alpha\approx 0.63$, and $n\approx 3.9$ for the $\bar{\nu}_{e}$ luminosity would suggest a SN explosion leading to a PNS of mass $1.44~{}M_{\odot}$, whose interior properties are described by the SFHo EoS. As further discussed in Sec. VII, due to the similarities of the spectra and luminosities of electron antineutrinos and heavy-lepton antineutrinos in our models at times $t\gtrsim 1$ s, flavor conversions are not a major effect, although they would make the reconstruction of the fit parameters less straightforward, but without spoiling our result. The accurate reconstruction of the parameters is left for future work.
@@ IV.2 LS220 EoS
  1. (i)
    The “normalization parameter” $C$ (top panels) is $C\sim O(10-1000)\,{\rm B}/\operatorname{s}$ for all of the neutrino species and it increases linearly with the PNS mass.
  2. (ii)
    The “power-law index” $\alpha$ (second panels from top) tends to decrease as the PNS mass increases. It is negative for all models and neutrino species except for $\nu_{e}$ and $\bar{\nu}_{e}$ in the simulation with $M_{\rm NS}=1.36\,M_{\odot}$, where $\alpha>0$.
  3. (iii)
    The “late-time suppression parameters” $\tau$ and $n$ (third and fourth panels from top) tend to decrease with higher PNS mass. For the 1.36-LS220 model we get $\tau\sim 4$ s and $n>1$ for $\nu_{e}$ and $\bar{\nu}_{e}$ (see Table 5 in Appendix B), whereas $\tau\lesssim 0.25$ s and $n\lesssim 0.5$ in all the other cases (Tables 6 and 7).
@@
Figure 11: Correlations of the fitting parameters $\tau$ (in units of seconds) and $\alpha$ for $\nu_{e}$ (top left panel), $\bar{\nu}_{e}$ (top right panel), $\nu_{\mu}$ (central left panel), $\bar{\nu}_{\mu}$ (central right panel), $\nu_{\tau}$ (bottom left panel), $\bar{\nu}_{\tau}$ (bottom right panel), for all Class A EoSs and all considered PNS masses, obtained from simulation output data between 1 s and $t_{\nu_{i},{\rm c}}$ for each neutrino species $\nu_{i}$.
@@ V Correlation between $\tau$ and $\alpha$
for all neutrino species except $\bar{\nu}_{e}$. This is shown in Fig. 11 for neutrinos (left) and antineutrinos (right) of all flavors and all Class A EoSs. In Appendix D, in Table 9 we report the best-fit values with errors obtained from linear regression for $A$ and $B$ for all neutrino species. The pairs of values $(\tau,\alpha)$ are represented by filled circles for the DD2 EoS, filled squares for SFHo, and filled diamonds for SFHx in Fig. 11. Different colors correspond to the different PNS masses: black for $M_{\rm NS}=1.36~{}M_{\odot}$, red for $M_{\rm NS}=1.44~{}M_{\odot}$, blue for $M_{\rm NS}=1.62~{}M_{\odot}$, green for $M_{\rm NS}=1.77~{}M_{\odot}$, and orange for $M_{\rm NS}=1.93~{}M_{\odot}$.
@@ VI Impact of muons and convection
In order to assess the impact of some of our physics inputs of the simulations, we investigate additional PNS cooling calculations for a PNS mass of 1.62 $M_{\odot}$ now, where we either omitted convection or muons. Non-convective models were computed for all considered EoSs and are denoted by a suffix “-c” appended to their names (e.g., 1.62-DD2-c), whereas only two models are considered without muons (suffix “-m”), namely 1.62-DD2-m and 1.62-SFHo-m [49]. The bottom two data blocks of Table 1 provide the final simulation times and the corresponding reduction factors $X_{\nu_{i}}^{\rm fin}$ for all neutrino species in these additional models. Bold numbers in Table 1 for models 1.62-SFHx-c and 1.61-LS220-c signal that the simulations were stopped when $X_{\nu_{i}}^{\rm fin}>0.15$, implying that for these models our standard cutoff time $t_{\nu_{i},{\rm c}}$ is larger than $t_{\rm fin}$. In both of the simulations, the weakest suppression is obtained for $\nu_{\mu}$, with $X_{\nu_{\mu}}^{\rm fin}=0.219$ for 1.62-SFHx-c and $X_{\nu_{\mu}}^{\rm fin}=0.185$ for 1.61-LS220-c (see the values marked by a star in Table 1). Therefore, to test the impact of convection for the SFHx and LS220 EoSs we cut our luminosity data at $t_{\nu_{i},{\rm Co}}$, when $X_{\nu_{i}}^{{\rm Co}}=t_{\nu_{i},{\rm Co}}\,L_{\nu_{i},{\rm Co}}/t_{\nu_{i},{\rm max}}\,L_{\nu_{i},{\rm max}}=0.22$ for simulations with SFHx and $X_{\nu_{i}}^{{\rm Co}}=0.19$ for simulations with LS220. The values of $t_{\nu_{i},{\rm Co}}$ for all neutrino species of these simulations are given in Table 4 of Appendix A.
@@
In the absence of convection the quantity $t\,L_{\nu}$ declines steeply only at $t\gtrsim 10\,\operatorname{s}$ for all neutrino species, with different characteristic features depending on the EoS (see the second data block from the bottom of Tables 2 and 3 in Appendix A). In particular, we witness the following:
  1. (i)
@@
  4. (iv)
    LS220 (red lines) leads to a peak in $t\,L_{\nu}$ at $t\approx 2$ s and shows a more shallow decline afterward, forming a plateau-like shape in the time interval $2~{}\operatorname{s}\lesssim t\lesssim 10$ s (i.e., $L_{\nu}$ follows approximately $L_{\nu}\propto t^{-1}$) before a steeper decline sets at $t\gtrsim 15$ s.
@@
  1. (i)
    the “absence of convection” leads to a considerable stretching of the PNS Kelvin-Helmholtz neutrino cooling time, with the most moderate change for DD2,
  2. (ii)
    the “omission of muons” has a relatively mild effect on the evolution of the neutrino signals for the displayed 1.62 $M_{\odot}$ models, making the suppression in the luminosity only slightly faster (because the NS becomes less compact with a lower binding energy), as visible by the green lines in the left panels.
To quantitatively assess the impact of these variations of the input physics of our models, we also fit the neutrino and antineutrino signals of the additional simulations with the expression of Eq. (2) and compare the best-fit parameters with those obtained in our benchmark simulations. As an example, Fig. 14 presents the simulation data (blue) and their best fits (orange) in the time interval of interest for $\bar{\nu}_{e}$ results from the simulations without convection (DD2 and SFHo in the top panels and SFHx and LS220 in the middle panels) and from our two simulations without muons (bottom panels). For the simulations without muons, where the neutrino signal does not experience major changes, the agreement between data and fits is similarly good as for our benchmark simulations. For the simulations without convection, the fit is still of excellent quality for DD2, although the omission of convection has altered the shape of the curve of $tL_{\bar{\nu}_{e}}$ (see Fig. 13). In contrast, we obtain visibly larger discrepancies between fits and data for the non-convective simulations with SFHo and SFHx, for which, in particular, the shape of $tL_{\bar{\nu}_{e}}$ in the power-law dominated early phase cannot be reproduced as well as for models that include convection. Notably, the fits for $\bar{\nu}_{e}$ (and similarly for all other neutrino species) slightly overestimate $tL_{\nu}$ at $t\approx 1\,\operatorname{s}$ and tend to peak only at somewhat later times. Finally, the plateau-like region of $tL_{\nu}$ in the simulation results for LS220 implies a best-fit value of $\alpha\approx 1$, with the luminosity fit overestimating the data at $t\approx 1\,\operatorname{s}$ and following well a $t^{-1}$ power law before being exponentially suppressed at $t\gtrsim 10\,\operatorname{s}$. Our findings for all other neutrino species are analogous. In Tables 10 and 11 of Appendix E we provide the best-fit parameter values and their $1\sigma$ errors for electron and muon neutrinos and antineutrinos, for our $1.62\,M_{\odot}$ models with different EoSs and varied input physics.
@@
which implies that the exponential luminosity decline starts at later times and also proceeds more slowly. For $\bar{\nu}_{e}$ in model 1.62-DD2-c we notice an exception from the described general trends with respect to $\tau_{\bar{\nu}_{e}}$, which is slightly smaller than the value of the corresponding model with convection (see Table 10 and Fig. 16 in Appendix E). In this case the mild decrease of $\tau_{\bar{\nu}_{e}}$ seems to be compensated by a reduction of $n_{\bar{\nu}_{e}}$ by a factor $\sim$3 compared to the non-convective model, which is by far the largest relative change for any neutrino species in all models with vs. without convection. Interestingly, the change in $\alpha$ depends on the EoS and neutrino species, showing, for instance, a decrease in 1.62-DD2-c compared to 1.62-DD2, an increase in 1.62-SFHo-c compared to 1.62-SFHo, and a decrease or slight increase in 1.62-SFHx-c compared to 1.62-SFHx depending on the type of neutrino. This nonuniform behavior points to differences in the influence of the EoS on PNS convection and the associated effects on the emission of different kinds of neutrinos during the early PNS cooling phase.
For simulations with the LS220 EoS, partly because of the poorer quality of the fits for the benchmark models, the omission of convection leads to radical changes in the values of the best-fit parameters. Indeed, LS220 simulations without convection show positive values of $\alpha$ (around unity), much larger values of $\tau$ and $n$ compared to the full-physics cases, and values of $C$ that are well compatible with those of simulations with the other EoSs including and excluding convection, i.e., the $C$ values are close to the luminosity values at 1 s instead of being several 100 B/s for our benchmark models. Finally, the weaker impact of muons on the neutrino signal is highlighted by the small changes in the best-fit parameters obtained for simulations without muons, as further detailed in Appendix E.
VII Counting rate in neutrino detectors
In order to exemplify a possible application of our luminosity fits, we discuss in this section the time evolution of the counting rate $tR_{\nu}(t)$ in a neutrino detector that will monitor the $tL_{\nu}(t)$ evolution in the case of a future Galactic SN explosion. We will demonstrate that our fitting recipe is also useful for fitting the observed neutrino signal. For this purpose, we consider as a reference case a SN at a distance of $D=10$ kpc and evaluate the predicted signal in the water Cherenkov detector of Super-Kamiokande (SK) [55, 56], inspired by the analysis in Ref. [57].